find the slopes, if they exist, of the linecontaining these points?
(-5,4)and(1-2)

(3,2)and(-5,0)

Answer by fractalier(6550)

Slope m is found by the formula
m = (y2 - y1) / (x2 - x1)
So for (-5,4) and (1,-2), we have
m = (-2 - 4) / (1 -(-5)) =
m = -6 / 6 = -1
And for (3,2) and (-5,0), we have
m = (0 - 2) / (-5 - 3) =
m = -2 / -8
m = 1/4